OpenStudy (anonymous):

Help!?!?!? 12. A sound wave is modeled with the equation y = 1/4 cos 2pi/3 θ. a. Find the period. Explain your method. b. Find the amplitude. Explain your method.

OpenStudy (anonymous):

@amistre64 @ganeshie8

OpenStudy (anonymous):

@.Sam. @jhonyy9 @RadEn

OpenStudy (anonymous):

@whpalmer4 @Hero

OpenStudy (whpalmer4):

\[y = \frac{1}{4}\cos(\frac{2\pi}{3}\theta)\] You know the period of the \(\cos\) function is \(2\pi\), right? Every \(2\pi\) that the argument to \(\cos\) goes through is one complete cycle. You've multiplied the argument by \(\frac{2\pi}{3}\) so that changes the period. If we had simply multiplied by 2, you would have 2 cycles per \(2\pi\) change in \(\theta\) instead of the usual 1, and the period would be correspondingly halved. Can you figure out the new period in your equation? The amplitude is easy to find. The value of \(\cos\) varies from 1 to -1. You've multiplied it by 1/4. What will that do to the amplitude?

OpenStudy (anonymous):

Im sorry, im not sure what you just said... Lol I really suck at math. I mean i REALLY suck. Can you show me step by step?

OpenStudy (anonymous):

Just a forewarning, it might look like i logged out, but im still going to be on and watching and commenting. :) Just didnt want you to be upset. :)

OpenStudy (anonymous):

You can keep helping tho!

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):